RC Meeting and Run Plan, Friday, February 15

From HallCWiki
Jump to: navigation, search

Main | Online | RC | Run Plan | Official Website | Redmine | Index | Nav

RC Meeting Information

RC Daily Meetings are at 4:30 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  • Jack, Ethan, Steve and Joe hooked up HMS Q2 power supply up to the magnet. It reached 600 A (we need 513A for first setting). Then need a long soak at some high current setting (which we will do while we run for now).
  • Ran C, LD2, Al dummy, LH2 targets for x>1 last night. Almost there, just needs to run other nuclear targets in the future.
  • DAQ needs to be swapped back to coincidence.
  • Need to think about optics/verification of Q2 at each setting (maybe off a couple amps....)

  • Summary of runs from last night:

1) x>1 LD2, Ep = -9.8, theta = 13.1 deg: run 7197, 7198, 7202-7205, 7207-7209

2) x>1 LH2, Ep = -9.8, theta = 13.1 deg: run 7210

3) x>1 Al, Ep=-9.8, theta = 13.1 deg: run 7201

4) x>1 1.5% C, Ep=-9.8, theta = 13.1 deg: runs 7199, 7200, 7206

Current Run Plan

  • Backup run plan: F2 positron running
  • Keep ELCLEAN rate <5 kHz
  • TO will need to adjust the cryo load between LH2 and LD2!
  • Use the F2 tune to set the magnets: go_magnets_SHMS_F2

1) 33 deg, Ep = +1.7 GeV , elclean statistics = 10 K on each target: LH2, LD2, Al dummy

2) 33 deg, Ep = +2.6 GeV , elclean statistics = 8 K on each target: LH2, LD2, Al dummy

3) 25 deg, Ep = + 3 Gev, elclean statistics = 3 K on each target: LH2, LD2, Al dummy

  • If there is time, take more statistics!

  • Sylvester's negative half of J/psi [1]

--> run at the last block of kinematics under "Setting scan" - 15 min per kinematic setting

Notices in Effect for Shift Crew

  • DO NOT RAMP HMS Q2 to 0!!!!!!!!! If it reaches 10A, just turn it off. Let the magnet settle for about 5 minutes at the set current before changing it.
  • We have no control over polarity for HMS magnets now. They don't have meaning in the GUI, either. It is currently set for positron/proton running (Q1-, Q2+ on hall probes)
  • If rotation cannot be Reset, call the RC, then call Ops for an escorted/controlled access.
  • Shift Checklists need to be done for every shift!

Current Issues

Planned Accesses to the Hall

  • Restricted Access all day to work on HMS Q2 power supply.

Planned Accelerator Configuration Changes

Special Requests